solve the equation: sinx+(√3/2)=0
Answer by stanbon(75887)

solve the equation:
sin(x)+(√3/2) = 0
-------
sin(x) = -sqrt(3)/2
------
Reference angle is 60 degrees.
But sine is negative in QIII and QIV::
So x = 180+60 = 240 degrees
Or x = 360-60 = 300 degrees
--------------
Cheers,
Stan H.
================